### v1

# RoboTransfer: Geometry-Consistent Video Diffusion for Robotic Visual Policy Transfer

Abstract
Imitation Learning has become a fundamental approach in robotic manipulation. However, collecting large-scale real-world robot demonstrations is prohibitively expensive. Simulators offer a cost-effective alternative, but the sim-to-real gap make it extremely challenging to scale. Therefore, we introduce RoboTransfer, a diffusion-based video generation framework for robotic data synthesis. Unlike previous methods, RoboTransfer integrates multi-view geometry with explicit control over scene components, such as background and object attributes. By incorporating cross-view feature interactions and global depth/normal conditions, RoboTransfer ensures geometry consistency across views. This framework allows fine-grained control, including background edits and object swaps. Experiments demonstrate that RoboTransfer is capable of generating multi-view videos with enhanced geometric consistency and visual fidelity. In addition, policies trained on the data generated by RoboTransfer achieve a 33.3% relative improvement in the success rate in the Diff-Obj setting and a substantial 251% relative improvement in the more challenging Diff-All scenario. Explore more demos on our project page: https://horizonrobotics.github.io/robot_lab/robotransfer
**footnotetext: Equal contribution. footnotetext: Corresponding author.
1 Introduction
Imitation Learning (IL) has become a fundamental approach for visuomotor control in robotic manipulation [67, 68]. However, collecting large-scale real-world robot demonstrations is prohibitively expensive [6, 7]. Simulated environments offer a cost-effective alternative [54, 59, 19, 44], but the scarcity of assets and sim-to-real gap[51, 43] make it extremely challenging to scale.
Recently, diffusion models [26] have gained attention as a promising method to synthetically generate realistic and diverse data. To expand robotic datasets without the need for extensive real-world data collection, ROSIE [62] employs text-guided image generation models trained on real-world data. However, it performs image augmentation on a per-frame basis, which leads to a loss of temporal and spatial consistency. In contrast, Cosmos-Transfer [2] generates photorealistic data using a video diffusion model conditioned on segmentation and depth, which helps preserve geometric consistency.
Despite the rapid advancements in the field, two key challenges remain. First, robotic systems often rely on observations from a single viewpoint, which inherently limits their perception. To overcome this limitation, multi-view observation is a commonly used approach. However, generating multi-view consistent results remains a significant challenge for video generative models. Second, robot manipulation scenarios are typically complex and highly interactive. Controlling such scenarios precisely through textual input alone, as demonstrated in text-to-video frameworks [5, 70], presents substantial difficulties.
In this work, we introduce RoboTransfer, as is shown in Figure 1, a video generation framework designed for data synthesis in robotics. To the best of our knowledge, it is the first framework to integrate multi-view geometry while offering explicit control over background, object attributes. RoboTransfer incorporates cross-view feature interactions and global depth/normal conditions within the video diffusion architecture, ensuring 3D consistency across different camera views and over time. To enhance controllability, RoboTransfer decomposes each scene into distinct control conditions, enabling background edits and object swaps by conditioning on this decomposition.
[…]
Figure 1: Overview of RoboTransfer.
Our work makes the following contributions:
  1. 1.
    We propose RoboTransfer, a robot generation framework based on multi-view, geometry-aware video diffusion models that maintain fidelity across multiple reference conditions.
  2. 2.
    We introduce a dataset construction pipeline that decomposes real-world recording data into high-quality triplets, which incorporate global geometry and appearance control conditions.
  3. 3.
    Through carefully designed experiments and evaluations, we demonstrate that our approach generates multi-view videos with significantly improved geometric consistency and visual fidelity. Furthermore, policies trained with RoboTransfer-generated data exhibit better generalization to novel and unseen environments.
2 Related Work
2.1 Visual Generalizable Imitation Learning
Imitation Learning (IL) has emerged as a foundational approach for visuomotor control in robotic manipulation. Deep networks trained on collected demonstrations—e.g., DP [12], OpenVLA [31], Act [67], Aloha [68], E2E-VP [36], and RT2 [7]—map raw visual inputs directly to motor commands. To bolster policy robustness and generalization, researchers have explored novel architectures [12] and vision pretraining schemes [4, 13]. However, gathering large-scale, task-specific robot data remains prohibitively expensive. Consequently, many methods leverage auxiliary sources—human demonstration videos [20], high-fidelity simulators [54, 59, 19, 44], or operational logs from other platforms [46, 30]. Each of these domains exhibits its own biases and visual characteristics, introducing distributional shifts that can degrade IL policy performance when transferred to the target environment. In this work, we offer a complementary perspective on generalizable visual imitation learning by leveraging state-of-the-art video diffusion models. By generating target-conditioned training data, our approach significantly enhances the robustness of policy and enables seamless transfer to novel environments.
2.2 Data Augmentation for Policy Learning
Data augmentation is widely used to train robust machine learning models. Prior work has applied augmentation directly to real-world robot captures to improve the resilience of visual policies against environmental variations[41, 42, 15, 23, 21]. Image-inpainting methods further enhance appearance invariance by replacing or masking out scene elements in collected frames [62, 11, 10, 40]. Likewise, neural view synthesis techniques generate novel viewpoints from existing robot data, promoting view-invariant feature learning [53, 8]. Simulators can cheaply produce vast quantities of labeled frames [54, 59, 19, 44], but discrepancies in physics modeling, rendering fidelity, and scene composition often hinder direct sim-to-real transfer. Domain randomization is used to close the sim-to-real gap, thereby enabling policies trained in simulation to generalize more effectively to real-world environments. [35, 22, 34, 51, 1, 48, 25]. In contrast to these methods, we leverage a video diffusion model to generate photorealistic image manipulations. This approach preserves both multi-view and geometry consistency, producing richly detailed samples for robot policy training.
2.3 Generation Models in Embodied AI
With the rapid advancement of generative models such as HunyuanVideo [33], Wan [56], SVD [5], Cosmos [2], CogVideoX [61], OpenSora [70], there is a growing trend toward leveraging generated data to bridge the gap between synthetic environments and real-world physical applications [72]. In particular, domains like autonomous driving have seen increasing adoption of generative techniques, as demonstrated by [57, 66, 27, 58, 17, 18, 64, 45, 65]. Similarly, robotic scenarios—where real-world data collection is often prohibitively expensive and time-consuming—have greatly benefited from generative data. Models such as UniSim [60], UniPi [14], and RoboDreamer [71] synthesize future robot behaviors via text prompts, which are then translated into actionable commands using inverse dynamics models. However, purely video-based generation methods often suffer from spatial and temporal inconsistencies, leading to degraded performance in downstream action prediction and planning. To address this, TesserAct [69] proposes a multi-modal generation pipeline that simultaneously synthesizes RGB, depth, and normal videos, enabling the reconstruction of coherent 4D scenes. This framework ensures both spatial and temporal consistency in robotic environments and supports policy learning that significantly surpasses prior video-only world models. Moreover, data generalization remains a critical challenge in robot learning. Techniques like ReBot [16] and RoboEngine [63] adopt background inpainting to diversify environmental textures, while Cosmos-Transfer [2] utilizes video-to-video translation to enrich overall scene appearance. Despite their success, these approaches lack fine-grained control over foreground and background textures, often resulting in synthetic data distributions misaligned with real-world tasks, especially when training policy models. Additionally, their inability to ensure multi-view consistency limits their scalability and effectiveness in complex, robotic settings.
3 Methods
We introduce RoboTransfer, a framework designed for the controllable generation of multi-view videos to support the training of policy models. By providing explicit control over both scene geometry and appearance, RoboTransfer enables the synthesis of video data with precisely defined distributions. This fine-grained control allows for the generation of diverse training scenarios, which are critical for improving the generalization of policy models. In the remainder of this section, we first present the preliminaries of video diffusion models in Sec. 3.1. Then, we describe the framework of RoboTransfer in Sec. 3.2. Finally, the dataset construction pipeline is elaborated in Sec. 3.2.
3.1 Video Diffusion Model Preliminaries
Diffusion-based methods for controllable video generation typically model the process as a gradual refinement of a noise latent $\boldsymbol{\epsilon}$ into a clean video latent $\mathbf{x_{0}}$, under the guidance of spatially aligned conditions $y_{s}$ (e.g., depth map, segmentation map) and unstructured conditions $y_{u}$ (e.g., CLIP embedding). Popular approaches in this paradigm include Flow Matching [37], DDPM [26], and EDM [29]. In the case of EDM, the learning objective is formulated as:
$\mathcal{L}(D_{\theta},\sigma)=\mathbb{E}_{\mathbf{x_{0}},y_{s},y_{u}}\left[\left\|\mathbf{x_{0}}-D_{\theta}\left(\mathcal{E}(\mathbf{x_{0}}+\mathbf{n}),\tau_{u}(y_{u}),\tau_{s}(y_{s}),\sigma\right)\right\|_{2}^{2}\right].$ (1)
Here, $\mathbf{x_{0}}\sim p_{\text{data}}$ denotes a clean sample drawn from the dataset, and $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ represents Gaussian noise. $\mathcal{E}$ denotes the encoder component of a VAE [32], and the denoising network $D_{\theta}$ is conditioned on the noise level $\sigma$ and encoded conditions $\tau_{u}(y_{u}),\tau_{s}(y_{s})$. To encourage stable learning across varying noise magnitudes, the total training objective aggregates the per-noise loss via a weighted expectation:
$\mathcal{L}(D_{\theta})$ $=\mathbb{E}_{\sigma}\left[\frac{\lambda(\sigma)}{\exp(u(\sigma))}\mathcal{L}(D_{\theta},\sigma)+u(\sigma)\right],$ (2)
$\lambda(\sigma)$ $=\frac{\sigma^{2}+\sigma_{\text{data}}^{2}}{(\sigma\cdot\sigma_{\text{data}})^{2}},$ (3)
$\ln(\sigma)$ $\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2}).$ (4)
In this setup, $\sigma_{\text{data}}$ denotes the empirical standard deviation of the data, while the distribution of $\sigma$ is governed by the hyperparameters $P_{\text{mean}}$ and $P_{\text{std}}$. The weighting function $\lambda(\sigma)$ ensures that all noise levels contribute proportionally during the initial stages of training.
[…]
Figure 2: The RoboTransfer framework performs multi-view consistent modeling to jointly reason across viewpoints. It represents geometry with scaled depth and normal maps, and encodes appearance using reference backgrounds and object-specific images for detailed control over appearance.
3.2 RoboTransfer Framework
The overall framework of RoboTransfer is illustrated in Figure 2. To ensure multi-view consistency during generation, we perform multi-view consistent modeling, enabling the generation process to reason jointly over information from different viewpoints. On the conditions side, RoboTransfer incorporates fine-grained control by encoding both geometric and appearance information. Specifically, we represent geometry using scaled depth maps and surface normal maps, capturing the underlying 3D structure of the scene. Meanwhile, the appearance is encoded using reference background images and object-specific images, providing detailed control over texture, color, and contextual appearance. In the following sections, we first introduce the multi-view consistent modeling, and then describe the encoding mechanisms for geometric and appearance conditions, respectively.
Multi-view Consistent Modeling. In robotic environments, it is common to deploy multiple cameras from different viewpoints to capture a scene simultaneously. To generate multi-view consistent videos for downstream tasks, RoboTransfer leverages the multi-view in-context learning capabilities [66, 28] inherently present in pretrained diffusion models. Specifically, given $N$ synchronized videos $\{V_{1},V_{2},...,V_{N}\}$ from different viewpoints, we concatenate them along the width dimension and encode the joint information using a VAE encoder $\mathcal{E}$:
$\mathbf{x}_{0}=\mathcal{E}([V_{1},V_{2},\ldots,V_{N}]),$ (5)
where $[V_{1},V_{2},\ldots,V_{N}]$ represents the concatenated multi-view video sequence. This modeling approach maintains the integrity of the pretrained video diffusion architecture, requiring no structural modifications, while enabling fast convergence and high-quality, view-consistent video generation in robotic settings.
Geometry Conditions. Recent video generation models primarily learn spatiotemporal coherence through data-driven approaches such as temporal attention [5, 70] or full attention [56, 33, 2]. While effective at capturing pixel-level dynamics, these methods lack an explicit understanding of underlying 3D geometry, limiting their applicability in robotic environments. To address this limitation, RoboTransfer explicitly incorporates geometry conditions to enhance the model’s awareness of scene structure and depth continuity. Specifically, we utilize depth and surface normal videos as geometric conditions. These two types of geometric cues are complementary — depth maps provide information about spatial distances from the camera, while normal maps encode local surface orientations. When combined, they offer a richer and more holistic description of scene geometry (see Sec. 3.2 for details on data acquisition). Since both the depth and normal videos are spatially aligned with the RGB views, we propose GuideNet, which utilizes a stack of convolutional layers that jointly downsample and encode the depth and normal sequences. The resulting geometry-aware representation is then concatenated with the noise latents along the channel dimension, allowing the generation process to be guided by accurate 3D structural cues.
Appearance Conditions. To enable fine-grained control over the generated textures, RoboTransfer introduces texture conditions from two complementary perspectives: background appearance and object appearance. Specifically, we use a background reference image and a set of object reference images as conditions signals to guide the generation process toward faithful reproduction of scene textures and object-level details. A key challenge in appearance control is ensuring compatibility with the geometry conditions. Naïvely combining appearance and geometry inputs may introduce conflicts (e.g., mismatched depth and texture), thereby weakening both the structural consistency and the visual fidelity of generated results. To address this, we carefully curate the appearance reference images, resulting in a background reference image $C_{b}$ and object reference images $C_{o}$ that do not contradict the geometric priors (see Sec. 3.2 for details). For background appearance, we employ a VAE encoder to compress $C_{b}$ into a latent representation that is spatially aligned with the generation latents. This encoded background appearance is then concatenated with the latent inputs to control the global texture and background style of the output video. In contrast, object appearance presents additional challenges due to the variable number and spatial distribution of objects in the scene. Therefore, object images are treated as unstructured conditions. We encode each object’s appearance using CLIP [47], which produces a global embedding used to guide the generation via cross attention. This design allows flexible and scalable conditioning on diverse object appearances while preserving compatibility with the structured geometry inputs.
[…]
Figure 3: RoboTransfer data processing pipeline. The data processing pipeline consists of two main components: Geometry conditions (left) are derived by using Mono-Normal and Video Depth Anything, which are scale-aligned with sensor depth to ensure consistency. Appearance conditions (right) are obtained by sampling keyframes from the video. GPT-4 is used to generate object descriptions, which are then processed by Grounding-SAM to create per-object masks. Additionally, background inpainting is used to generate complete reference backgrounds.
3.3 Dataset Construction from Real World Data
Data construction is central to our framework. The overall pipeline is illustrated in Figure 3. Unlike conventional text-to-video or image-to-video tasks, our approach requires both globally consistent geometry and multiple part-specific reference images (e.g., background and manipulable objects).
Our pipeline begins with the collection of a large-scale robotic demonstration dataset. Each demonstration captures synchronized RGB-D streams from both a hand-eye and an overhead RGB-D camera. To cover a wide range of operating conditions, we vary backgrounds (e.g., textured tabletops, cluttered surfaces) and manipulable objects (shapes, sizes, materials).
To enforce spatial and temporal consistency, we incorporate geometric cues such as depth and surface normals. Since raw depth maps from RGB-D sensors are often noisy, we employ the state-of-the-art video depth estimator, Video Depth Anything[9], to produce temporally consistent depth maps. We then align the scale of the estimated depth map with the RGB-D sensor’s depth using robust least-squares fitting to ensure global spatial accuracy. Additional details are provided in the supplementary material. For surface normals, we compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS[24], which jointly serves as reliable geometric conditions.
Next, we sample a subset of RGB keyframes as visual references. These frames are processed by GPT-4 to generate structured scene descriptions, including object-level descriptions for individual entities. These descriptions guide the Grounding-SAM module[50], which detects and segments per-object masks. To obtain clean background references, we inpaint the segmented objects from the original images, generating plausible object-free scenes (e.g., an empty tabletop). For multimodal object-level conditions, the segmented object masks are further processed with the CLIP model[47] to obtain semantic embeddings. These conditions allow the diffusion model to generate photorealistic, semantically grounded video augmentations that respect both object identity and spatial layout.
4 Experiments
4.1 Synthesis Quality
Implementation Details. To construct our training dataset, we leverage the open-source Cobot Magic platform to collect a large-scale video corpus of dual-arm robot executions, covering diverse tasks such as grasping and placing. The raw videos are segmented into 10Hz clips of 30 frames each, resulting in approximately 24k clips for training. Additionally, we curate a set of 1.6k 10Hz 30-frame clips from the collected dataset for video synthesis quality evaluation. All videos are annotated with conditions as described in Sec. 3.2, to facilitate both training and evaluation. RoboTransfer is fine-tuned from the pretrained video diffusion model [5]. During training, videos from each camera view are resized to a resolution of $640\times 384$. We adopt the AdamW optimizer with a learning rate of $3\times 10^{-5}$ and a global batch size of 8, training for a total of 70K steps. During inference, we use the EDM scheduler [29] to perform 30 denoising steps and apply classifier-free guidance.
Table 1: Experiment comparison on different geometry conditions of RoboTransfer.
Model Camera RMSE $\downarrow$ Abs.Rel. $\downarrow$ Sq.Rel. $\downarrow$ Mean Err. $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
RoboTransfer [D.S.] left 0.074 0.124 0.020 4.86 2.88 142.90 218.51
RoboTransfer [D.P.] left 0.054 0.090 0.010 3.91 2.28 149.68 123.31
RoboTransfer [Scaled D.P.] left 0.049 0.081 0.008 3.48 1.99 183.26 112.43
RoboTransfer [Scaled D.P. + N.] left 0.047 0.079 0.008 3.31 1.92 202.03 107.43
RoboTransfer [D.S.] middle 0.182 0.074 0.031 3.51 1.58 153.76
RoboTransfer [D.P.] middle 0.132 0.053 0.015 3.05 1.43 95.89
RoboTransfer [Scaled D.P.] middle 0.134 0.054 0.015 2.93 1.40 103.32
RoboTransfer [Scaled D.P. + N.] middle 0.133 0.054 0.015 2.86 1.39 101.17
RoboTransfer [D.S.] right 0.090 0.137 0.025 4.93 2.91 40.70 396.33
RoboTransfer [D.P.] right 0.072 0.103 0.016 4.14 2.36 56.02 262.96
RoboTransfer [Scaled D.P.] right 0.064 0.090 0.011 3.74 2.11 65.45 226.76
RoboTransfer [Scaled D.P. + N.] right 0.058 0.087 0.009 3.55 2.00 75.67 220.12
Evaluation Metrics. We evaluate the generated videos from multiple perspectives, including multi-view consistency, geometric consistency, and semantic consistency. For multi-view consistency, we follow [3] and utilize the state-of-the-art image matching method [52] to compute the number of matched pixels (Mat.Pix.) between adjacent views (i.e., left-to-center and right-to-center). For geometric consistency, we adopt the evaluation framework from Cosmos-Transfer [2], which assesses depth and surface normal alignment. For depth prediction, we compute scale-invariant metrics including Root Mean Squared Error (RMSE), Absolute Relative Error (Abs.Rel.), and Squared Relative Error (Sq.Rel.). For normal estimation, we report the Mean Angular Error (Mean Err.) and Median Angular Error (Med. Err.). In terms of semantic consistency, we measure appearance controllability using CLIP [47] similarity. For background-level similarity, we compute the CLIP cosine similarity between a reference background image and each generated frame (BG. Sim.). For foreground object similarity, we use GroundingDINO [38] and SAM2 [49] to segment objects in each generated frame and compute their CLIP similarity with a reference image (Obj. Sim.). Due to motion and occlusion in the wrist-mounted (left and right) views, object tracking becomes unreliable. Therefore, we evaluate this metric only on the center view to ensure accuracy. All the above metrics are computed on a per-frame basis, and we report the average value across all frames as the final score. Additionally, we compute the FVD [55] between the generated and real videos to further assess the overall realism and temporal coherence of the generated data.
Geometry Consistency Analysis. We first validate the geometric consistency in RoboTransfer’s generative capability. As shown in Table 1, directly using raw depth sensor data (D.S.) as a conditional input introduces considerable noise, which adversely affects the generation quality. This degradation is evident across all three views in terms of depth accuracy, surface normals, and multi-view consistency. In contrast, employing model-predicted [9] depth (D.P.) results in smoother depth maps, which serve as a more stable supervisory signal for training the generative model. Consequently, the generated videos exhibit significantly better geometric consistency. Specifically, compared to D.S.-based generation, D.P.-based conditions relatively improves the RMSE and Mean Err. of the middle view by 27.4% and 14.2%, and increases Pix.Mat. of the middle and left/right views by 4.7% and 37.6%, respectively. Further improvements are observed where we incorporate scaled predicted depths (Scaled D.P.) to enforce a consistent depth scale across multiple views. This multi-view scale alignment further enhances geometric consistency, leading to relative improvements of 22.4% and 16.8% in Pix.Mat. of left and right views. Finally, combining normal maps with depth as joint conditions (Scaled D.P.+N.) provides complementary geometric cues, yielding the best overall performance across all geometric metrics.
Appearance Consistency Analysis.
Table 2: Experiment comparison on different appearance conditions of RoboTransfer.
Bg. Inpaint Obj. Split Camera BG. Sim. $\uparrow$ Obj. Sim. $\uparrow$ RMSE $\downarrow$ Med.Err. $\downarrow$ Pix.Mat. $\uparrow$ FVD $\downarrow$
left 0.796 0.051 1.94 191.59 117.25
left 0.802 0.049 1.92 198.32 119.65
left 0.797 0.050 1.93 196.85 118.22
left 0.805 0.047 1.92 202.03 107.43
middle 0.712 0.847 0.137 1.46 108.65
middle 0.719 0.845 0.135 1.41 105.66
middle 0.712 0.855 0.136 1.42 105.54
middle 0.720 0.858 0.133 1.39 101.17
right 0.753 0.064 2.12 71.71 226.84
right 0.762 0.060 2.04 72.44 223.91
right 0.754 0.061 2.05 72.68 222.11
right 0.764 0.059 2.00 75.67 220.12
We then evaluate RoboTransfer’s capability in controlling visual appearance. As shown in Table 2 (row-1 and row-2 of all three camera views), we observe that directly conditioning on background images without inpainting can corrupt the geometric cues. In contrast, inpainting the background before conditioning not only preserves the geometric structure but also improves appearance consistency, leading to $\sim 1\%$ relative improvements in RMSE, median error, and background similarity. Comparing row-1 and row-3 of the middle camera view, we examine object-level appearance control. In row-1, the entire image with all the objects is encoded using a global CLIP feature. In contrast, row-3 applies a finer-grained approach by masking and encoding each object individually, followed by feature concatenation. This object-wise conditions leads to a 1% improvement in object CLIP similarity, indicating better control over individual object appearances. Finally, row-4 of all three camera views demonstrate that combining background inpainting with object-wise CLIP encoding achieves the best performance in both foreground and background appearance control. This setup yields the highest CLIP similarity scores for both regions, confirming the effectiveness of jointly modeling structured and unstructured visual components.
Visualizations. We demonstrate that RoboTransfer enables controllable manipulation of both foreground and background appearance while preserving geometric and multi-view consistency across camera views. As shown in Figure 4, RoboTransfer successfully modifies the background appearance while retaining the foreground object’s texture and scene geometry. Figure 5 illustrates that the model can alter the foreground appearance without affecting the background or spatial layout. Finally, Figure 6 demonstrates the model’s ability to jointly edit both foreground and background elements. All visualizations confirm that RoboTransfer maintains geometric and multi-view consistency throughout the generated scenes.
[…]
Figure 4: Visualizations of RoboTransfer with different background reference images.
[…]
Figure 5: Visualizations of RoboTransfer with different object reference images.
[…]
Figure 6: Visualizations of RoboTransfer with different background and object reference images.
4.2 Real Robot Experiments
To evaluate RoboTransfer, we conducted experiments on a real-world robotic platform, assessing how synthetic data impacts visual policy generalization across varied environments.
Experimental Platform and Setup Details. Our experiments were conducted on the Agilex Cobot Magic platform, equipped with two PIPER robotic arms and three Intel RealSense D435i cameras—two mounted on the wrists and one positioned overhead. We used only RGB data for policy learning, creating a challenging visual understanding problem while reflecting realistic deployment constraints.
Benchmark Task and Evaluation Methodology. We selected a dual-arm manipulation sequence with four phases: (i) left arm grasps a spoon; (ii) spoon is placed on a tray; (iii) right arm grasps the spoon; (iv) spoon is placed in a cup. We evaluated across three test conditions: identical objects/background (Same), novel object appearances (Diff-Obj), and variations in both objects and environment (Diff-All). Testing protocols are detailed in the supplementary material. We introduced a Stage Score (0-4) corresponding to successfully completed phases, providing finer-grained insights than binary success/failure metrics. This reveals how far policies progress through the task sequence, particularly valuable when analyzing generalization.
Implementation Details. We adopted ACT [67] as our baseline architecture without modifications. For each task, we collected 100 expert demonstrations via ALOHA [68] teleoperation. Our methodology follows a pretrain+finetune paradigm. We trained baseline policies on real data for 100k steps, then fine-tuned with synthetic data for 50k additional steps (learning rate reduced to 1/10). For the grasp spoon task, training on real data required 8 NVIDIA H20 GPUs for approximately 24 hours, while the synthetic data fine-tuning phase utilized the same hardware configuration for about 12 hours. Additional details of the training protocol are given in the supplementary material.
Effectiveness of Synthetic Data Augmentation.
Table 3: Performance comparison across different test conditions and models. Success Rate (%) and Stage Score (0-4) are reported for each configuration. The best results for each test conditions are highlighted in bold.
Model Same Diff-Obj Diff-All
Success (%) Score Success (%) Score Success (%) Score
Baseline (Real only) 80.0 3.6 33.3 2.7 13.3 1.6
+ Obj Augmentation 44.4 3.0 22.2 2.0
+ Obj+Bg Augmentation 46.7 3.0
Table 3 presents the comparative results. Results show significant performance degradation when baseline policies face novel visual contexts, with success rates dropping from 80.0% to 13.3%. However, policies fine-tuned with our generated data show considerable resilience. In Diff-Obj conditions, success rates improved by $33.3\%$ (relative), while in Diff-All scenarios, we observed a $251\%$ relative improvement. Stage Score results further confirm that augmented models make more progress through task phases, indicating more robust manipulation understanding even when tasks aren’t fully completed.
Effect of Synthetic Data Proportions. We investigated how different proportions of synthetic data affect performance by evaluating synthetic ratios of 25%, 50%, 75%, and 100%. Figure 7 shows these results for the Diff-All conditions. Performance follows a non-linear relationship with synthetic data proportion, peaking at 50% and showing diminishing returns beyond. Success rates increased from 13.3% with no synthetic data to 46.7% at 50% synthetic ratio—a 3.5x improvement. Stage scores similarly peaked at 3.0 with 50% mixture versus 1.6 without synthetic data. Notably, using exclusively synthetic data (100%) yielded 40.0% success rate and 2.9 stage score—improvements over purely real data but below the balanced 50% mixture. This confirms complementary contributions: synthetic data provides diversity while real data contributes physical fidelity, with optimal interaction at equal proportions. RoboTransfer preserves geometric consistency while enabling systematic appearance variation, addressing a fundamental challenge in robot learning. Additional details on model configuration and extended evaluations are in the supplementary material.
Figure 7: Performance metrics across synthetic data mixing ratios. Left: Success rate vs. ratio. Right: Average stage score vs. ratio. Note that 0% represents our baseline model trained only on real data. Both metrics peak at 50% synthetic data, suggesting an optimal balance between synthetic diversity and real-world physical fidelity.
5 Conclusion
In this work, we proposed RoboTransfer, a novel diffusion-based data synthesis framework in robotics, which integrates multi-view geometry while offering explicit control over background and object attributes. We also introduced our dataset processing pipeline, which provides triplets incorporating both global geometry and appearance control conditions. The experiments and evaluations show that our RoboTransfer can produce multi-view consistency data for training robotic manipulation policies. Therefore, the generalization of visuomotor policy gets significantly improved.
Limitation. Although the method is general and flexible, several limitations remain that we aim to address in future work. First, our approach currently focuses on augmenting the appearance of objects and scenes, rather than generating new motions. Moving forward, we plan to integrate RoboTransfer with additional simulators to broaden its applicability. Furthermore, we aim to scale up both the size and diversity of the training data, which will enhance the variety and robustness of the generated outputs.
Appendix
[…]
1
Appendix A Dataset Construction Details
For the training of RoboTransfer, we collected data and designed a dedicated dataset construction pipeline (see Figure 3). In this section, we provide additional implementation details.
A.1 Data Collection
Robot Platform: We construct a large-scale robotic demonstration dataset using the Agilex Cobot Magic platform, adhering to standardized data collection protocols established in prior work [44, 39]. The robot is equipped with two PIPER robotic arms. As illustrated in Figure 8, each demonstration includes synchronized RGB-D streams captured from three Intel RealSense D435i cameras: one mounted at two hand-eye viewpoint and one positioned overhead.
Dataset Design: To capture a broad spectrum of manipulation scenarios, we systematically vary tasks, backgrounds, and manipulable objects. Specifically, we design twelve distinct manipulation tasks (visualized in Figure 9), each representing diverse interaction patterns. For each task, we collect 100 demonstration segments across 10 unique object configurations, resulting in a total of 1,000 samples per task. The dataset incorporates a variety of backgrounds, including textured tabletops and cluttered surfaces, to reflect realistic operating conditions. Manipulable objects are diversified in terms of shape, size, and material, enhancing both the visual richness and physical complexity of the dataset. This comprehensive design ensures a high degree of diversity and realism, which is crucial for training robust and generalizable visual policies.
[…]
Figure 8: Visualizations of robot platform.
[…]
Figure 9: Visualizations of the dataset.
A.2 Geometry conditions
Depth conditions: To enforce both spatial and temporal consistency, we incorporate geometry-conditioned supervision into our dataset construction pipeline. During data collection, we obtain synchronized RGB-D streams, including multi-view depth maps. However, depth maps from the RGB-D sensors are often noisy and incomplete, particularly in regions with low reflectivity (e.g., black surfaces), leading to domain gaps across samples. In contrast, depth maps rendered from simulators are typically complete and noise-free.
While we explored existing depth completion methods, we found that many suffer from limited temporal consistency and poor generalization across diverse scenes. To address these limitations, we adopt the state-of-the-art Video Depth Anything (VDA) model [9], which produces temporally coherent and spatially complete depth maps.
Despite the high-quality temporal consistency of depth estimates from Video Depth Anything (VDA), these predictions lack a globally consistent scale across multiple views. To address this, we align the VDA-predicted depth maps with the RGB-D sensor measurements via a robust least-squares scale fitting procedure. While standard least-squares regression offers a closed-form solution for affine scale alignment, it is susceptible to errors introduced by sensor noise and depth outliers. To mitigate this, we introduce a robust multi-frame fitting strategy (Figure 10) that aggregates depth correspondences across time and iteratively filters outliers based on residual error thresholds. This iterative refinement ensures stable and accurate recovery of the global scale transformation. Our approach effectively preserves the geometric completeness provided by VDA while anchoring depth predictions to the real-world scale captured by the RGB-D sensor. The resulting alignment ensures both spatial accuracy across views and temporal smoothness across frames, enabling reliable depth conditions.
Input:
Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
Output:
Scaled Depths $\mathbf{D}_{\text{scaled}}\in\mathbb{R}^{B\times H\times W}$,
1
2Initialize Mask: $\mathcal{M}\leftarrow(\mathbf{D}_{\text{sensor}}>\epsilon)\land(\mathbf{D}_{\text{pred}}>\epsilon)$ ;
3
4for 2 iterations do
5       $\mathcal{E}\leftarrow|\mathbf{D}_{\text{pred}}-\mathbf{D}_{\text{sensor}}|\odot\mathcal{M}$ ;
6       $\tau\leftarrow\mathcal{P}_{80}(\mathcal{E}[\mathcal{M}>0])$ ;
7       $\mathcal{M}\leftarrow(\mathcal{E}<\tau)\odot\mathcal{M}$ ;
8       $s,b\leftarrow\textbf{Scale Fitting}(\mathbf{D}_{\text{pred}},\mathbf{D}_{\text{sensor}},\mathcal{M})$ ;
9       $\mathbf{D}_{\text{scaled}}\leftarrow s\cdot\mathbf{D}_{\text{pred}}+b$ ;
10
return $\mathbf{D}_{\text{scaled}}$;
Algorithm 1 Dynamic Mask Refinement
Input:
Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
Mask $\mathcal{M}\in\{0,1\}^{B\times H\times W}$
Output: Scale $s\in\mathbb{R}$, Shift $b\in\mathbb{R}$
1 Flatten and Mask:
$\mathbf{p}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{pred}}\odot\mathcal{M})\in\mathbb{R}^{N}$
$\mathbf{s}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{sensor}}\odot\mathcal{M})$
[…]
2Construct Global System:
$\begin{bmatrix}\mathbf{p}^{\top}\mathbf{p}&\mathbf{p}^{\top}\mathbf{1}\\ \mathbf{p}^{\top}\mathbf{1}&\mathbf{1}^{\top}\mathbf{1}\end{bmatrix}\begin{bmatrix}s\\ b\end{bmatrix}=\begin{bmatrix}\mathbf{p}^{\top}\mathbf{s}\\ \mathbf{1}^{\top}\mathbf{s}\end{bmatrix}$
Closed-Form Solution:
$s=\frac{(\mathbf{1}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{1}^{\top}\mathbf{s})}{\det}$
$b=\frac{(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})}{\det}$
where $\det=(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{1})-(\mathbf{p}^{\top}\mathbf{1})^{2}>\epsilon$
Algorithm 2 Scale Fitting
Figure 10: Depth Scale Alignment: Implementation Details
Normal conditions: For high-precision surface reconstruction, surface normals are critical and can provide surface details rather than only depth. At the same time, normal does not have scale consistency issues. We compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS [24].
A.3 Appearance conditions
Object Mask: For the manipulation task data, we first select keyframes that effectively capture both the objects and the background conditions of the scene. We then extract all operable objects using a structured prompt template (see Figure 11), which describes each object in terms of its color, material, shape, and spatial position. These object descriptions are fed into Grounding DINO[38] to obtain bounding boxes for object detection. The resulting bounding boxes are subsequently used as prompts for SAM2[49] to generate the corresponding object masks.
*System Instruction*
As a visual descriptor, articulate objects with this flow:
Color -> Material -> Shape/Texture -> Spatial Position
*Template Format*
"A [color] [material] [object], [shape] with [texture], located [region](x,y)"
*Demonstration Examples*
- A red plastic ball, spherical with a smooth glossy surface, at the center (250,300)
- A brown wooden chair, angular with a matte wood grain, in top-left (100,50)
- A silver-blue metallic arm, mechanical with a brushed texture, on the right (600,200)
Figure 11: Visual Description Prompt Template Architecture.
Objects and Background conditions:To extract the object conditions, we first segment image patches corresponding to individual objects from the full image. Each patch is resized to 384×384 and passed through the CLIP model [47] to obtain RGB embeddings. For the background conditions, we mask out all detected objects from the full image using their corresponding masks. The masked image is then fed into an inpainting model to reconstruct the background with the objects removed.
Appendix B Visual Policy Implementation Details
To validate the effectiveness of the data generated by RoboTransfer, we train a visual policy using the procedures detailed below for both training and deployment.
B.1 Data Collection and Preprocessing
To ensure the validity of our evaluation, we excluded the dataset described in Section A from the real-robot experiments. The data preparation proceeded as follows:
Real Expert Data: We collected 100 sets of expert demonstration data for each manipulation task using the ALOHA teleoperation system. The primary observations were RGB images (1280x720 resolution, subsequently downscaled to 640$\times$360 to optimize training), with camera streams captured at approximately 30Hz and sampled at 10Hz for policy training. Auxiliary robot state information (e.g., joint positions, end-effector poses) was originally recorded at 200Hz and downsampled to 50Hz for model use.
Synthetic Data Generation: To further enhance model performance and generalization, we generated synthetic data based on the collected real data, specifically targeting variations in foreground objects and background scenes, for use in the subsequent fine-tuning phase.
B.2 Training Pipeline and Core Parameters
Our training pipeline is based on the ACT (Action Chunking Transformer) architecture, which processes visual input from three distinct perspectives (two wrist-mounted cameras and one fixed overhead camera). At each training timestep, the model receives a single RGB frame from each view.
Training Objective and Two-Stage Training Strategy: The policy model is designed to predict a sequence of the next 100 robot states, corresponding to a 2-second action horizon given the 50Hz robot state frequency.
We adopted a "pre-train-then-finetune" paradigm:
Pretraining on Real Data:The model was first pretrained for 100k steps using the collected real expert demonstration data. During this phase, the batch size was set to 512, and the learning rate was $1\times 10^{-4}$.
Finetuning with Synthetic Data: After pretraining, the aforementioned synthetic data was introduced to fine-tune the model for an additional 50k steps. The learning rate for this phase was reduced to $1\times 10^{-5}$.
All training was performed on a cluster equipped with 8 NVIDIA H20 GPUs. The pretraining phase took approximately 24 hours, and the fine-tuning phase required about 12 hours.
B.3 Real-Robot Deployment and Evaluation
The trained policies were deployed and evaluated in the following environment:
Deployment Platform: Experiments were conducted on the Agilex Cobot Magic platform8, the same platform used for initial data collection, ensuring consistency between training and testing environments.
Inference Procedure: During deployment, the policy operated in a synchronous execution mode. Specifically, the system waits for the robot to fully execute the action sequence (comprising 100 actions) generated from the model’s previous inference step before capturing new observations and proceeding to the next model inference and subsequent execution. At each decision point, the model receives as input a single RGB frame from each of its multiple camera views and the current robot state and outputs the sequence of 100 actions. The latency for this single model inference process is approximately 10ms. These actions are then executed on the robot at a 50Hz control frequency, consistent with the robot state prediction rate during training.
Appendix C More Qualitative Results
C.1 Real-to-Real Transfer
The proposed RoboTransfer offers strong controllability over both foreground and background generation. Given the same structured input, the model allows for flexible editing of background attributes such as texture and color, as shown in Figure 12. Meanwhile, as illustrated in Figure 13, the appearance of foreground objects, including their color, can also be effectively modified. This real-to-real generation framework enriches the diversity of training data for downstream tasks, contributing to better generalization of the policy model.
[…]
Figure 12: Visualizations of RoboTransfer with different background reference images.
[…]
Figure 13: Visualizations of RoboTransfer with different object reference images.
C.2 Sim-to-Real Transfer
As shown in Figure 14, RoboTransfer generates photorealistic videos from simulated structural inputs, including out-of-distribution cases. This Sim-to-Real paradigm reduces dependency on structured information from real-world datasets, thereby better supporting downstream robotic learning tasks.
[…]
Figure 14: Visualizations of Sim-to-Real generation results.
C.3 Comparison
Figure 15 illustrates that Cosmos Transfer [2] is mainly effective for robot scene generation under fixed camera settings (central view), while its quality drops notably for dynamic camera viewpoints (side views). Compared to Cosmos Transfer [2], RoboTransfer demonstrates significantly better multi-view consistency, enabling more realistic and coherent novel-view synthesis across varying camera poses.
[…]
Figure 15: Comparison with Cosmos generation results.